OpenStudy (anonymous):

Simplify tan(arcsin x +arccos X) for 0 =< x =< 1 (=< means less than or equal to)

OpenStudy (anonymous):

Let\[\theta =(\sin^{-1} x +\cos^{-1} x)\]Make a right triangle with angle tan theta.